Is H3PO4 a triprotic or diprotic acid, if it is a trprotic acid, then how do we explain its reaction with sodium hydroxide? i.e. H3PO4 + 2NaOH == Na2HPO4 + 2H2O.
1
Expert's answer
2012-04-12T09:53:05-0400
H3PO4 is a trprotic acid,but the reactivity of cleavage ща proton determine the dissociation constant. For the third proton is much smaller than for the second and first
H3PO4(s) + H2O(l) https://upload.wikimedia.org/wikipedia/commons/thumb/9/96/Equilibrium.svg/15px-Equilibrium.svg.png H3O+(aq) + H2PO4−(aq)
Ka1 = 7.25×10−3
H2PO4−(aq) + H2O(l) https://upload.wikimedia.org/wikipedia/commons/thumb/9/96/Equilibrium.svg/15px-Equilibrium.svg.png H3O+(aq) + HPO42−(aq)
Ka2 = 6.31×10−8
HPO42−(aq) + H2O(l) https://upload.wikimedia.org/wikipedia/commons/thumb/9/96/Equilibrium.svg/15px-Equilibrium.svg.png H3O+(aq) + PO43−(aq)
Ka3 = 3.98×10−13